X squared plus 10x =24
Answer by shree840(260)

x^2+10x=24
x^2+10x-24=0
x^2+12x-2x-24=0
x(x+12)-2(x+12)=0
(x-2)(x+12)=0
either x-2=0 or x=2
or x+12=0 x=-12 ANS